# SPECIAL RELATIVITY

### v9

@@ 7 Notion of Ordinary Space in Special Relativity
We introduce new approach to the aberration of light theory in non-inertial frames of reference, finding another way in which this complicated problem can be solved. For example, when viewed from the inertial frame, the aberration of light effect in the accelerated frame (Fig. 16) is easy to calculate in the framework of the special theory of relativity taking advantage of the relativity of simultaneity. In contrast, in order to compute the aberration of light effect in non-inertial systems we used a metric tensor.
@@ 7.1 Inertial Frame View of Observations of the Noninertial Observer
First we want to rise the following interesting and important point. The laws of physics in any inertial reference frame should be able to account for all physical phenomena, including the observations made by non-inertial observers. It is widely believed that all phenomena in non-inertial (e.g. rotating) reference system should be considered only in the framework of the space-time geometric approach using Langevin metric. However, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well known relativistic kinematic effects.
@@
For example, when viewed from the inertial lab system, the interpretation of the Sagnac effect is simple and the phase difference (attributed to the Sagnac effect) between counter-propagating waves may be derived from the relativistic law of velocity composition. In other words, Sagnac effect in the rotating frame, as viewed from the inertial lab frame, presents a kinematic effect (Einstein’s velocity addition) of special theory of relativity. It is implicitly assumed that the speed of light is $c$ in the lab frame which, in accordance with Einstein special relativity, is independent of motion of the source. In contrast, the Sagnac effect is not easy to calculate in a frame of reference attending rotation. In this case, authors of textbook used a metric tensor (Langevin metric) in a plane four-dimensional Minkowski space-time to calculate the propagation time difference between counter-running waves. Can we not look at the aberration of light effect in the same way? Here also we used the Langevin metric in a non-inertial frame of reference.
@@
Let us first consider the application of the classical kinematic method to the computation of the aberration of light effect in a non-inertial frame, as viewed from an inertial frame. Authors of textbooks obtain the same addition theorem for light beam velocities as for particle velocities in the classical physics. This result is particularly remarkable because the study of light aberration is intimately connected with old kinematics. If we neglect $v^{2}/c^{2}$ comparison with unity, the Galilean vectorial law of addition of velocities is actually used. However, the classical kinematics method leads to a serious mistake if applied to the computation of the aberration of light effect. Let us consider the aberration of light in the inertial frame. According to our approach, there is a remarkable prediction concerning the deviation of the energy transport for light transmitted through a hole on a moving screen, Fig. 9. In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric. The crossed term in this metric generates anisotropy in the accelerated frame that is responsible for the change of radiation direction, Fig. 14. In contrast, based only on the (Galilean or Einstein’s) velocity addition, we arrive at the conclusion that there is no aberration of the transmitted light beam in the accelerated frame.
@@
The error in the last argument follows from the fact that incorrect assumption that an inertial observer and accelerated observer have common 3-space is used. According to the special relativity, inertial observer and accelerated (with respect to the fixed stars) observer have different 3-spaces. Errors inherent in the classical kinematics method applied to the computation of the aberration of light effect are due to the use of the velocity of light simply as a classical velocity. Relativistic effects do not have a place in this description. According to classical approach, there is no principal difference between the aberration of light and the aberration of raindrops. We first notice that one of the postulate of special relativity states that if two distinct events cannot be connected by a causal signal that travels no faster than a light signal, then they cannot be connected by a causal signal at all. We will show that the aberration of light effect is a corollary to the relativistic kinematic effects. The appearance of relativistic effects in radiation phenomena does not depends on a large speed of the radiation sources. Light is always a relativistic object. In particular the relativity of simultaneity is responsible for aberrations to the first order in $v/c$.
@@
Now, it is very interesting to show that the geometric effects in our ordinary space world are closely associated with the relativity of simultaneity. In Fig. 16 the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light. The question cannot be avoided relative to what a light beam propagated in the accelerated frame with angular displacement $-v/c$? Suppose that an observer in the accelerated frame performs the direction of the light beam measurement and the plane wavefront of transmitted light beam is imaged by a lens to a diffraction spot which lies in the focal plane on the optical axis. Measurement of the direction of the optical axis with respect to the frame axes is equivalent to the determination the angular displacement. In order to detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed. We must inquire in detail by what method we assign coordinates. This method involves some sort of physical procedure; eventually it must be such that it will be give us coordinates in both (inertial and accelerated) frame of reference.
Figure 32: Aberration of light in the accelerated frame of reference. In the Einstein time synchronization for the screen, the accelerated observer describes the transmitted light beam and the reference beam from an emitter at rest using usual Maxwell’s equations. The $x_{n}$-axis is parallel to the wavefront of the reference beam. The aberration increment may be defined as an angle between the reference and transmitted light beams.
Figure 33: Aberration of light in the inertial frame of reference. In the absolute time synchronization for the screen, the inertial observer describes the reference beam from an emitter at rest using usual Maxwell’s equations. The $x$-axis is parallel to the wavefront of the reference beam. The aberration increment may be defined as an angle between the reference and transmitted light beams.
@@
In ordinary space we find that the accelerated frame moves with respect to the inertial frame along the line motion and the inertial frame moves with respect to the accelerated frame along the line motion. The angle between the axis of the observer’s coordinate system and the line motion is a simple ordinary space geometric parameter. Using the line motion as a reference $x_{n}$-axis, the accelerated observer can then define the second reference axis. We need to give a ”practical”, ”operational” answer to the question of how to assign an axis perpendicular to the $x_{n}$-axis. Clearly, it is possible to define a reference direction using a light beam. We will define the second reference direction in the following way. Let us suppose that the aberration direction inside the accelerated frame is determined with reference to the fixed direction of light beam from a ”plane-wave” emitter which is at rest in the accelerated frame. In other words, a coordinate system is formed here by electromagnetic axis and line motion. The motion of the aberrated light beam are assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
@@
When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all direction. In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations. According to Maxwell’s electrodynamics, light is always emitted in the direction normal to the radiation wavefront. We consider the case in which the velocity component of the (reference) light beam along the $x_{n}$-axis is equal to zero. In other words, the $x_{n}$-axis is parallel to the wavefront of the reference ”plane wave”, Fig. 32. The number that specifies the aberration increment of the light beam transmitted through the hole may be defined as an angle between the reference and transmitted light beams.
Figure 34: Inertial frame view of observations of the accelerated observer. In the Lorentz coordinatization, the reference light beam in the accelerated frame is emitted in the direction normal to the wavefront. If make a Lorentz boost, we introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just an inclination of the electromagnetic axis in the accelerated frame.
@@
Let us try to get an understanding of the relationship between the reference directions inside the inertial frame and the reference directions inside the accelerated frame. Such approach in our case of (aberration of light) interest is based on the use of local (reference) light sources. We determine the reference directions perpendicular to the line motion in the inertial frame and the accelerated frame using two local light sources. In effect, one source should be stay at rest in the accelerated frame (Fig. 32) while the second should be at rest in the inertial frame (Fig. 33). In other words, the reference electromagnetic axis in each frame is formed by an individual light beam. Of course, it is possible to find the space description on a more practical level than that of light beams. For instance, the reference axis in the earth-based frame may be formed by the gravitation field vector (e.g. a standard direction of plumb-line). The equivalence of all local physical frames of reference underlies the theory of relativity, so when the aberration angle emerges, it emerges in the same manner in all local physical reference systems. A comparison with a light clock might help here. When inertial observer looks at the light clock inside the accelerated frame, he sees that clock run slowly. Not only does this particular kind of clock run more slowly, but if the theory of relativity is correct, any other clock, operating on any principle whatsoever, would also appear to run slower.
@@
Now let us return to observations of an accelerated observer, as viewed from an inertial frame. Let us start with non-covariant approach. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. According to our approach, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through a hole, Fig. 9. However, there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists in using Lorentz coordinatization to describe the uniform translation motion of the light source in the inertial frame. In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with rule clock structure in the inertial frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. Now speed of light is independent of motion of the source in the inertial frame. This means that the velocity of light relative the moving system is $c-v$ in the positive $x$-direction, and $-c-v$ in the negative direction. It is immediately understood that, as a result of motion of accelerated system along the $x$-axis (i.e. along the plane of simultaneity of the reference beam in the accelerated frame) with velocity $v$, there is relativity of simultaneity. This reasoning is in analogy with Einstein’s train-embankment experiment. On the one hand, the wave equation remains invariant with respect to Lorentz transformations.On the other hand, if use the Lorentz coordinatization , we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the accelerated frame on the angle $v/c$.
@@
From what has preceded it is easy to see that in the accelerated system, from the point of view of the inertial observer, the electromagnetic axis, assigned by the accelerated observer, will not parallel to the $z$-axis of the inertial observer, Fig. 34. We arrive at the conclusion that in the accelerated frame transmitted light beam propagates along the $z$-axis of the inertial frame of reference with angular displacement $-v/c$ with respect to the reference beam (i.e. with respect to the $z_{n}$-axis) in the accelerated frame.
@@
To continue our discussion of inertial frame view of observations of the non-inertial observer, let us consider the opposite situation. We determine the reference directions in the inertial frame and the accelerated frame using two local light sources. This has interesting consequences. For example, suppose that an observer in the inertial frame performs the direction of transmitted light measurement relative to the reference light beam from the emitter which is at rest with respect to the inertial frame (Fig. 33). The inertial observer would find that the angular displacement is positive and equal to $\theta_{a}=v/c$.
@@
Now we ask about the angular displacement of transmitted light with respect to the reference beam inside the accelerated frame. On the one hand, we know that angular displacement of the transmitted light with respect to the reference light beam from the emitter which is at rest with respect to the accelerated frame is $-v/c$. On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$, Fig. 17. We come to the conclusion that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the reference light beam of the inertial frame. In fact, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$. This result is consistent with the inertial observer measurements (Fig. 33), as must be.
@@ 7.2 The Non-Existence of an Instantaneous 3-Space
We derived the results for observations of the non-inertial observer with the help of Lorentz transformations. Let us assume, according to textbooks, that the line motion of observers is the same in the accelerated frame as in the inertial frame. At first site, if reference axes were orthogonal to the line motion they must be parallel to each other. Since there exist the angular displacement $v/c$, the situation seems paradoxical.
Figure 35: The lab observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the Wigner rotation.
First of all, as noted above, the key idea in resolution of this paradox is the realization of the fact that there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity we have a mixture - of positions and time. In other words, in the space measurement of one observer there is mixed a little bit of the time as seen by the other. The point, then, is not the existence of a common ordinary space. The fact that two observers with different trajectories have different 3-spaces is not considered in textbooks.
If we look at the situation carefully we see that the accepted in textbooks assumption of common (x) line motion of observers is based on the belief that observers have common 3-space. This is misconception. Right way would be to use the light beam as the common reference direction. We have already seen that the inertial observer can directly measure the angular displacement between the inertial ($z$) and accelerated ($z_{n}$) electromagnetic reference directions and this result is consistent with the accelerated observer measurement. In other words, the displacement between electromagnetic axes is a physical reality. Let us accept that the electromagnetic $z_{n}$ axis in the inertial frame is the reference light beam in the accelerated frame. Now the lab observer sees the rotation of the moving frame axes with respect to the lab frame axes, Fig. 35. This rotation is a relativistic kinematic effect. One can verify directly that the axes of accelerated frame are actually rotated with respect to the lab frame axes by the angle equal to $\theta_{w}=v/c$ in the first order in $v/c$.
Let us clear up the meaning of the parameter $\theta_{w}$. It can be associated with the angle of the so-called Wigner rotation (see the Section 7.5 for more details). Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of light beam observations of the non-inertial observer in perpendicular geometry. This inclination angle $v/c$ can be interpreted as a rotation angle $\theta_{w}=v/c$ of the moving axes with respect to the lab frame axes.

### v10

@@ 7.1 Inertial Frame View of Observations of the Noninertial Observer
We introduce a new approach to the aberration of light theory in non-inertial frames of reference, finding another way in which this complicated problem can be solved. First, we want to raise the following interesting and important points. The laws of physics in any inertial reference frame should be able to account for all physical phenomena, including the observations made by non-inertial observers. It is widely believed that all phenomena in non-inertial (e.g. rotating) reference systems should be considered only in the framework of the space-time geometric approach using the Langevin metric. However, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well-known relativistic kinematic effects.
@@
For example, when viewed from the inertial lab system, the interpretation of the Sagnac effect is simple, and the phase difference (attributed to the Sagnac effect) between counter-propagating waves may be derived from the relativistic law of velocity composition. In other words, the Sagnac effect in the rotating frame, as viewed from the inertial lab frame, presents a kinematic effect (Einstein’s velocity addition) of the special theory of relativity. It is implicitly assumed that the speed of light is $c$ in the lab frame which, in accordance with Einstein’s special relativity, is independent of the motion of the source. In contrast, the Sagnac effect is not easy to calculate in a frame of reference attending rotation. In this case, authors of textbooks used a metric tensor (Langevin metric) in a plane four-dimensional Minkowski space-time to calculate the propagation time difference between counter-running waves. Can we not look at the aberration of light effect in the same way? Here we also used the Langevin metric in a non-inertial frame of reference.
@@
Let us first consider the application of the classical kinematic method to the computation of the aberration of light effect in a non-inertial frame, as viewed from an inertial frame. Authors of textbooks obtain the same addition theorem for light beam velocities as for particle velocities in classical physics. This result is particularly remarkable because the study of light aberration is intimately connected with old kinematics. If we neglect $v^{2}/c^{2}$ comparison with unity, the Galilean vectorial law of addition of velocities is used. However, the classical kinematics method leads to a serious mistake if applied to the computation of the aberration of light effect. According to our approach, there is a remarkable prediction concerning the aberration of light phenomena in the accelerated frame, Fig. 20. To predict the result of the aberration measurement the accelerated observer should use the Langevin metric. The crossed term in this metric generates anisotropy in the accelerated frame that is responsible for the change of radiation direction. In contrast, based only on the (Galilean or Einstein’s) velocity addition, we conclude that acceleration does not spoil the symmetry between the accelerated reference frame and the initial inertial frame.
@@
The error in the last argument follows from the fact that the incorrect assumption that an inertial observer and accelerated observer have a common 3-space is used. According to the special relativity, inertial observer and accelerated (with respect to the fixed stars) observer have different 3-spaces. Errors inherent in the classical kinematics method applied to the computation of the aberration of light effect are due to the use of the velocity of light simply as a classical velocity. Relativistic effects do not have a place in this description. According to the classical approach, there is no principal difference between the aberration of light and the aberration of raindrops.
@@
Now, it is very interesting to show that the geometric effects in our ordinary space world are closely associated with the relativity of simultaneity. In Fig. 16 the transmitted light beam is propagated at the angle $-v/c$, yielding the phenomenon of the aberration of light. The question cannot be avoided relative to what a light beam propagated in the accelerated frame with angular displacement $-v/c$. Suppose that an observer in the accelerated frame performs the direction of the light beam measurement and the plane wavefront of the transmitted light beam is imaged by a lens to a diffraction spot that lies in the focal plane on the optical axis. Measurement of the direction of the optical axis with respect to the frame axes is equivalent to the determination of the angular displacement. To detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed. We must inquire in detail by what method we assign coordinates. This method involves some sort of physical procedure; eventually, it must be such that it will give us coordinates in both (inertial and accelerated) frames of reference.
@@
In ordinary space, we find that the accelerated frame moves with respect to the inertial frame along the line motion and the inertial frame moves with respect to the accelerated frame along the line motion. The angle between the axis of the observer’s coordinate system and the line motion is a simple ordinary space geometric parameter. Using the line motion as a reference $x_{n}$-axis, the accelerated observer can then define the second reference axis. We need to give a ”practical”, and ”operational” answer to the question of how to assign an axis perpendicular to the $x_{n}$-axis. It is possible to define a reference direction using a light beam. We will define the second reference direction in the following way. Let us suppose that the aberration direction inside the accelerated frame is determined concerning the fixed direction of the light beam from a ”plane-wave” emitter which is at rest in the accelerated frame. In other words, a coordinate system is formed here by electromagnetic axis and line motion. The motion of the aberrated light beam is assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle.
@@
When the accelerated system starts moving with constant velocity the standard procedure of Einstein’s clock synchronization can be performed. The Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. In this synchronization, the accelerated observer describes the light beam from an emitter at rest using usual Maxwell’s equations. According to Maxwell’s electrodynamics, light is always emitted in the direction normal to the radiation wavefront. We consider the case in which the velocity component of the (reference) light beam along the $x_{n}$-axis is equal to zero. In other words, the $x_{n}$-axis is parallel to the wavefront of the reference ”plane wave”. The number that specifies the aberration increment of the light beam transmitted through the hole may be defined as an angle between the reference and transmitted light beams.
Figure 30: Inertial frame view of observations of the accelerated observer.
@@
Let us try to get an understanding of the relationship between the reference directions inside the inertial frame and the reference directions inside the accelerated frame. Such an approach in our case of (aberration of light) interest is based on the use of local (reference) light sources. We determine the reference directions perpendicular to the line motion in the inertial frame and the accelerated frame using two local light sources. In effect, one source should stay at rest in the accelerated frame while the second should be at rest in the inertial frame. In other words, the reference electromagnetic axis in each frame is formed by an individual light beam. Of course, it is possible to find the space description on a more practical level than that of light beams. For instance, the reference axis in the earth-based frame may be formed by the gravitation field vector (e.g. a standard direction of plumb-line). The equivalence of all local physical frames of reference underlies the theory of relativity, so when the aberration angle emerges, it emerges in the same manner in all local physical reference systems.
@@
Now let us return to observations of an accelerated observer, as viewed from an inertial frame. Let us start with the non-covariant approach. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. According to our approach, when one has a transversely moving screen with a hole in it and a plane wave of light is falling normally on the screen, there is a deviation of the energy transport for light transmitted through a hole, Fig. 9. However, there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists in using Lorentz coordinatization to describe the uniform translation motion of the light source in the inertial frame. In Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with rule clock structure in the inertial frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. Now speed of light is independent of the motion of the source in the inertial frame. This means that the velocity of light relative to the moving system is $c-v$ in the positive $x$-direction, and $-c-v$ in the negative direction. It is immediately understood that, as a result of the motion of the accelerated system along the $x$-axis (i.e. along the plane of simultaneity of the reference beam in the accelerated frame) with velocity $v$, there is the relativity of simultaneity. This reasoning is in analogy with Einstein’s train-embankment experiment. On the one hand, the wave equation remains invariant with respect to Lorentz transformations. On the other hand, if use the Lorentz coordinatization, we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the accelerated frame on the angle $v/c$.
@@
From what has preceded it is easy to see that in the accelerated system, from the point of view of the inertial observer, the electromagnetic axis, assigned by the accelerated observer, will not parallel to the $z$-axis of the inertial observer, Fig. 30. We conclude that in the accelerated frame transmitted light beam propagates along the $z$-axis of the inertial frame of reference with angular displacement $-v/c$ with respect to the reference beam (i.e. with respect to the $z_{n}$-axis) in the accelerated frame.
@@
We determine the reference directions in the inertial frame and the accelerated frame using two local light sources. This has interesting consequences. First of all, the accelerated observer and the inertial observer can directly measure the angular displacement of the transmitted light beam with respect to the reference light beam of the inertial frame. Suppose that an observer in the inertial frame performs the direction of transmitted light measurement relative to the reference light beam from the emitter which is at rest with respect to the inertial frame. The inertial observer would find that the angular displacement is positive and equal to $\theta_{a}=v/c$.
@@
Now we ask about the angular displacement of transmitted light with respect to the reference beam inside the accelerated frame. On the one hand, we know that the angular displacement of transmitted light with respect to the reference light beam from the emitter which is at rest with respect to the accelerated frame is $-v/c$. On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$, Fig. 17. We conclude that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the reference light beam of the inertial frame. In fact, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$. This result is consistent with the inertial observer measurements, as must be 353535 We have been talking about the electromagnetic axis. Suppose that we determine the reference directions in the inertial frame and the accelerated frame using two relativistic electron sources. We emphasize an important difference between a reference light beam and a reference relativistic electron beam. In our experimental arrangement, it would be impossible to measure the angular displacement of transmitted light exactly using the reference electron beam axis..
@@ 7.2 The Non-Existence of an Instantaneous 3-Space
We derived the results for observations of the non-inertial observer with the help of Lorentz transformations. Let us assume, according to textbooks, that the line motion of observers is the same in the accelerated frame as in the inertial frame. At first sight, if reference axes were orthogonal to the line motion they must be parallel to each other. Since there exists the angular displacement $v/c$, the situation seems paradoxical.
Figure 31: The inertial observer sees the rotation of the moving axes with respect to the lab frame axes, which is closely associated with the relativity of simultaneity.
First of all, as noted above, the key idea in the resolution of this paradox is the realization of the fact that there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity, we have a mixture - of positions and time. In other words, in the space measurement of one observer, there is mixed a little bit of the time as seen by the other.
If we look at the situation carefully we see that the generally accepted assumption of common (x) line motion of observers is based on the belief that observers have a common 3-space. This is a misconception. The right way would be to use the light beam as a reference direction. We have already seen that the inertial observer can directly measure the angular displacement between the inertial ($z$) and accelerated ($z_{n}$) electromagnetic reference directions and this result is consistent with the accelerated observer measurement. In other words, the displacement between electromagnetic axes is a physical reality. Let us accept that the electromagnetic $z_{n}$ axis in the inertial frame is the reference light beam in the accelerated frame. Now the inertial observer sees the rotation of the moving frame axes with respect to the inertial frame axes, Fig. 31. This rotation is a relativistic kinematic effect. One can verify directly that the axes of the accelerated frame are rotated with respect to the inertial frame axes by the angle equal to $\theta_{w}=v/c$ in the first order in $v/c$ (see the next section for more details).